# MOND as a transformation between non-inertial reference frames via Sciama’s interpretation of Mach’s Principle

### v4

@@ 3 Discussion
It is shown that Milgrom’s MOND correction to Newtonian gravity or inertia can be reformulated based on Mach’s principle without dimensional constants or free parameters except for the speed of light, ensuring mass, space and time scale invariance at all regimes.
[…]
@@
The relationship pointed out by Milgrom between his MOND’s acceleration scale constant $a_{0}$ and the cosmological constant or the Hubble parameter is interpreted merely as accidental due to the cosmological coincidence by which, at the present epoch, the energy densities associated with dark energy (the cosmological constant) and visible matter are of the same orders of magnitude.
The Machian MOND’s transformation is a function $\mu(M_{u},R_{u},M,r)$ (4) between different scales and non-inertial reference frames: the relationship between a local frame in spacetime and that of a global frame which is at rest (the background of the rest of the universe) with respect to the first one, with both being accelerated relative to each other. It expresses how gravity or inertia changes locally due to the background universe through relative masses and distances in both systems, in accordance with relational mechanics, and respects the weak equivalence principle since it does not depend on the inertial mass. The speed of light in Sciama’s relationship takes the meaning of the speed of gravity and causality justifying the choice of $R_{u}$ as the radius of the observable universe containing mass $M_{u}$, which is the part of the universe causally affecting local dynamics, for Mach’s interpretation not to violate locality.
In Milgrom’s MOND, the external field effect, which is behind the non-linearity of MOND, does not allow to ignore the external field (the field of the mother system) when considering internal dynamics. But somehow the field of the universe, in which all systems and embedded into, is ignored in Milgrom’s MOND (if considered, all bodies in the universe would be above $GM_{u}/R_{u}^{2}$ and thus, in the high acceleration regime). By effectively replacing $a_{0}$ for $M_{u}/R_{u}^{2}$ in Machian MOND, the accelerated motion of the universe is considered due to the relativity of acceleration and inertia, when acceleration is defined with respect to the inertial frame of the universe à la Mach.
@@
Although the origin of the Machian MOND transformation is not explained or derived here, it seems that it follows from Mach’s principle’s argument about which frame is acceleration defined with respect to. When the local field intensity is above that of the rest of the universe (at the cores of galaxies or within the solar system), the ‘main frame’ acceleration is defined with respect to is the local one, and one can describe the dynamics of the system without referring to the field intensity of the universe at large through the usual Newtonian laws. But, if the field intensity of the rest of the universe is greater than the one from the local system (in the deep MOND regime, such as the flat velocity part of the disk in galaxies), it seems that acceleration is defined with respect to the background frame of the universe, and the field intensity of the universe at large must be taken into account non-linearly together with that of the local frame, to explain the dynamics of the local system (in the case of Machian MOND, effectively through the Machian MOND transformation (4) ). In between both regimes, the identification of the ‘main frame’ is shared between the local and global frames, described approximately by the slope of the Machian MOND transformation. This suggests that wherever local field intensities are higher than those of the universe at large, the Machian effects vanish (be that the one from Machian MOND, or from a varying gravitational constant in Brans-Dicke theory, in which tests showing the theory to be almost identical to GR are performed only at the high field intensity regime). And according to Mach, wherever local field intensities are below those of the universe at large, inertia is smaller and velocities are higher. This points towards the long-discussed problem of boundary conditions of the universe at large in GR, which has been linked to Mach’s principle in the past.
@@
Machian MOND satisfies several definitions of Mach’s principle: Newton’s gravitational constant $G$ is a dynamical field (and the inertial mass of a body increases with the agglomeration of masses in its neighborhood), local inertial frames are affected by the cosmic distribution of matter, and rotation is undefined up to the speed of light in the absence of the rest of the universe (or equivalently, a body in an empty universe has no inertia). In agreement with Mach’s principle, if inertial mass is defined by the potential of the mass distribution (following Sciama’s argument), it is reasonable to think that it could also be affected by the field intensity of the mass distribution. Thus, Machian MOND introduces an even stronger (and non-linear, due to the external field effect of MOND) dependence of local dynamics on the distribution of matter than linear theories that just consider Sciama’s relationship, and it can be considered more Machian.
Milgrom himself favors modifying inertia (which is the true aim of Mach’s principle), and not just gravity. An example of modified inertia is that of SR $F=m_{i}d(\gamma v)/dt$ with $\gamma$ being the Lorentz factor [37]. The original Milgrom’s MOND as modified inertia with the inertial term of the equation of motion $\mu\left(g/a_{0}\right)m_{i}g$ is translation, rotation, and Galilei invariant, but leads to non-conservation of momentum in a many-body systems because the equation of motion is not derivable from an action. As a modification of the kinetic term in the Lagrangian, it conserves momentum, energy, and angular momentum, but is not Galilei invariant. A non-relativistic MONDian kinetic action for a particle with conservation laws of momentum, energy, and angular momentum, and the boost symmetry being Galilean, must be time non-local [10]. But this only occurs if it is a function of the trajectory $r(t)$ (since the function requires knowledge of the entire trajectory) and it is required to have a MOND limit when $a_{0}\rightarrow 0$, which is neither the case in Machian MOND. We have omitted the original negative sign of Sciama’s relationship $c^{2}\sim-GM_{u}/R_{u}$ for simplicity. But by considering it and changing the sign in the Machian MOND transformation (4) from positive to negative for consistency, the resulting transformation
$\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)=\frac{1}{\sqrt{1-\left(\frac{-M_{u}/{R_{u}}^{2}}{M/r^{2}}\right)}}$ (5)
@@
resembles the structure of a Lorentz-like transformation $\gamma=1/\sqrt{1-\beta}$. If a function depending on velocities, such as the Lorentz factor, arises from Lorentz invariance in SR, where accelerations are absolute for consistency of the theory, one would expect a function depending on accelerations or field intensities, such as Machian MOND, to arise when relativizing accelerations. Lorentz invariance implies that the speed of light is constant in all inertial frames, and it seems that Machian MOND implies that the total field intensity at any point in the universe is always $>M_{u}/R_{u}^{2}$. It may be that Machian MOND arises from a new boost symmetry generalizing Lorentz invariance, for instance, through the relativity of acceleration, which is broken by the presence of the universe at large. Without a background universe, this symmetry would reduce to the Galilean or Lorentzian one. If inertia has an origin in gravitation, then the modification on the kinetic term can be based only on field intensities, without modifying the Poisson equation (and leaving gravity intact), which is shown in Machian MOND as modified inertia. The resulting Machian MOND formulation can be thought of as an in-between interpretation of modified inertia and modified gravity.
@@
Machian MOND, or a similar and equivalent approximate formulation, is considered to be a necessary non-linear feature that a phenomenological theory of modified inertia or modified gravity which incorporates Mach’s principle reduces to as an approximation, in order to agree with galaxy rotation dynamics. This might arise by imposing Mach’s principle (for instance, through the relativity of accelerations and inertia of local objects being determined by the sources of the gravitational field at the global scale) to a non-Machian non-linear Lagrangian or to a theory of gravity such as GR (in which solutions to local dynamics do not depend on the mass or size of the observable universe). In fact, analogous interpolating functions in SR and GR have, as independent variables, dimensionless ratios of the form $v/c$ and $MG/Rc^{2}$, similar to Machian MOND. The transformation function should arise naturally as an approximation, without the need of introducing it by hand (as in AQUAL or QUMOND), but the global $M_{u}$ and $R_{u}$ parameters governing local dynamics should appear in the complete FundaMOND theory, which is not the case in GR. So far, only linear and non-Lorentz invariant Machian theories of modified inertia with Sciama’s relationship have been constructed, such as those of Treder and Schrödinger.
@@
It is well known that, according to Bekenstein, a higher value for $a_{0}$ could resolve the tension between Milgrom’s MOND and galaxy cluster’s core dynamics (and that $a_{0}$ seems to grow with scale), where Newtonian accelerations take around that same value (provided that $a_{0}$ stays the same in galaxies for agreement with rotation curves). Machian MOND, although being equivalent to Milgrom’s MOND, is effectively a model of a varying $a_{0}$ with time scale and could in principle help resolving this issue of Milgrom’s MOND. It could also in principle solve the cosmological issues of Milgrom’s MOND [38], such as those arising from a fixed scale-dependence on $a_{0}$. As modified gravity, the equivalent variation of the Newtonian $G$ is both temporal due to the expanding universe and dependence on the mass and radius of the observable universe, and spatial since local potentials and local field intensities should also be included in Sciama’s relationship and the Machian MOND transformation, respectively.
@@
It would be interesting to compare Machian MOND with observational constraints of a varying effective gravitational constant at the early universe considering a cosmological model without most or all physical dark matter (assuming that Machian MOND solves at least partly the need for dark matter), due to its dependence on the mass and radius of the observable universe, which have both varied over time. A similar study has been done for quantized inertia by showing the prediction of a specific increase in the galaxy rotation anomaly at higher redshifts [39], but quantized inertia only makes the $a_{0}$ of Milgrom’s MOND dependent on $R_{u}$ through $a_{o}\sim c^{2}/R_{u}$. The formulation of quantized inertia is similar to the one proposed (without Sciama’s relationship), but its interpretation is not based on Mach’s principle. In fact, no Planck constant appears in the formulation, which suggests that its origin is not related to quantum phenomena, while the need to account for the universe at large directly points towards Mach’s principle. More recently, it has been observed that there is no significant evolution in the baryonic Tully-Fisher law for redshifts up to $z\sim 2$ [40], implying an almost constant $a_{o}$ in MOND. In contrast, Machian MOND shows a difference dependence compared to quantized inertia: the Machian MOND function depends not only on $R_{u}$, but also on $M_{u}$. If the term $M_{u}/R_{u}^{2}$ is more constant than the simple dependence on $R_{u}$, this observation would not rule out Machian MOND (the observable radius was smaller in the past, but so was the mass contained in the volume defined by the observable radius). But as explained before, Machian MOND depends on many uncertainties, such as the Hubble tension, whether some physical dark matter is present in $M_{u}$, missing integers and Pi in the formulation, and relativistic effects of the universe as a whole.
The form of the transformation (4) is not unique and depends on the chosen Milgrom’s MOND interpolating function for which the Machian substitutions are made. Although (4) might be in conflict with solar system constraints, as it approaches $\mu(x)\rightarrow 1$ too slowly at the high acceleration regime, this is a problem of the slope of the transformation, which can be chosen to be different by another form for it. The exact form of the transformation should not be fundamental but derived as an approximation from a complete FundaMOND theory of modified inertia á la Mach considering the relativity of acceleration and inertia. Furthermore, which integers and mathematical constants play a role within the Machian transformation and within Sciama’s relationship is left to be discussed for a better agreement with observations.
@@ Acknowledgments
Authors have no competing interests to declare. This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.

### v5

@@ 3 Discussion
Since only equivalent Machian substitutions for $G$ and $a_{0}$ are made yielding approximately the same numerical values (up to omitted numerical factors), and since $M$ and $r$ take the same meanings as in Milgrom’s MOND, the resulting corrections (3) and (4) for galaxy rotation curves are equivalent to Milgrom’s MOND at present epoch. For $\frac{M_{u}/{R_{u}}^{2}}{M/{r}^{2}}<<1$, $\mu\left(\frac{M_{u}/{R_{u}}^{2}}{M/{r}^{2}}\right)\rightarrow 1$ and Newtonian laws are restored with $v=c\sqrt{({M/r})/({M_{u}/R_{u}})}$. For $\frac{M_{u}/{R_{u}}^{2}}{M/{r}^{2}}>>1$, $\mu\left(\frac{M_{u}/{R_{u}}^{2}}{M/{r}^{2}}\right)\rightarrow\sqrt{\frac{M/r^{2}}{M_{u}/{R_{u}}^{2}}}$, $v=c\sqrt[4]{M/M_{u}}$, and asymptotic rotational velocities are determined by the relative gravitationally active masses of the nearby and large scale mass distributions.
This framework is free from dimensional constants and free parameters except for the speed of light, which takes the meaning of the speed of gravity and causality justifying the choice of $R_{u}$ as the radius of the observable universe containing mass $M_{u}$, which is the part of the universe causally affecting local dynamics.
The numerator $GM_{u}/{R_{u}}^{2}$ in (4) prior to canceling of the gravitational constants with the denominator $a_{N}=GM/r^{2}$ does not represent a gravitational field intensity, since these must vanish for large scale: at any point in the universe there exists a surrounding observable volume sphere of approximately constant matter density whose net field intensity is zero at its center. Two Machian possibilities for term $GM_{u}/R_{u}^{2}$ are left. It arises as either the average acceleration of masses in the universe radially receding from the observer, or as a directionless sum of the inverse-squared distance contributions from all masses in the observable universe, times the gravitational constant. The former would imply that for the case of a non-expanding universe, MOND effects would vanish and that the decrease of inertia is due to such radial motion of the distant masses. The latter, which seems more straightforward, implies that $a_{0}=3GM_{u}/R_{u}^{2}$ from the integral over the spherical observable volume. Such quantity will be referred to as the universal field scale, which is a scalar measure of how strongly mass is concentrated around an evaluation point. In Machian theories, the gravitational constant and inertial mass are sourced by a scalar field composed of inverse-distance mass contributions $\Phi=\sum_{i}m_{i}/|r-r_{i}|$, commonly identified with the Newtonian potential. Within this perspective, MOND can be interpreted as a nonlinear additional contribution arising from the scalar inverse-squared distance mass terms $\phi=\sum_{i}m_{i}/|r-r_{i}|^{2}$. This introduces a stronger dependence of local dynamics on the global matter distribution than linear approaches with only Sciama’s relationship, making the framework more explicitly Machian.
@@
The relationship pointed out by Milgrom between his MOND’s acceleration scale constant $a_{0}$ and the cosmological constant or the Hubble parameter is interpreted as merely accidental due to the cosmological coincidence by which, at the present epoch, the energy densities associated with dark energy (the cosmological constant) and visible matter are of the same order of magnitude.
@@
Although the origin of the Machian MOND function is not explained or derived here, it seems that it follows from Mach’s principle’s argument about which frame is acceleration defined with respect to. When local field intensities are above the universal field scale (at the cores of galaxies or within the solar system), the ‘main frame’ acceleration is defined with respect to is the local one, and one can describe the dynamics of the system without referring to the scalar field intensity of the universe at large through the usual Newtonian laws (and Machian effects vanish). But, if the scalar sum of field intensities of the rest of the universe is greater than the field intensity of the local system (in the deep MOND regime, such as the flat velocity part in galaxies), it seems that acceleration is defined with respect to the background frame of the universe, and the field scale of the universe at large must be taken into account non-linearly together with that of the local frame, to explain the dynamics of the local system (in the case of Machian MOND, effectively through (4) ). In between both regimes, the identification of the ‘main frame’ is shared between the local and global frames, described approximately by the slope of the Machian MOND function.
Curiously, while a decrease of inertia occurs the smaller the Newtonian field is with respect to the universal field scale when $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}>>1$, which agrees with the Machian notion of a decrease of inertia when (gravitational) masses are removed from the neighborhood of the body, a decrease of $M_{u}$ implies a vanishing of MOND effects and an increase of inertia ‘Newtonizing’ motion, which does not align with the Machian statement. However, by decreasing $M_{u}$ considering also Sciama’s relationship for the gravitational constant, as introduced in the observable $v=c\sqrt[4]{M/M_{u}}$, the gravitational force increases (having the same observational consequence as considering Sciama’s relationship alone when $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}<<1$, which is the decrease of inertia).
For the non-physical case in which the observable universe were infinite ($M_{u}\rightarrow\infty$, $a_{0}=GM_{u}/R_{u}^{2}\rightarrow\infty$), it corresponds to the deep MOND limit described pure MOND behavior in which the equations must be fully scale invariant. To achieve this, Milgrom sets $G\rightarrow 0$ keeping ${{\mathcal{A}}_{0}}\equiv Ga_{0}$ constant, arguing that Newtonian gravity introduces a dimensional constant $G$ with a preferred scale that breaks scale invariance. But in Machian MOND, $M_{u}\rightarrow\infty$ directly implies $G=c^{2}/(M_{u}/R_{u})\rightarrow 0$. This case would be the classical Newtonian cosmological model of an infinite, homogeneous, and isotropic universe, in which problems with the consistency of Newton’s laws have for long been discussed. Relatively recently, it has been argued that the solution comes from the introduction of a sort of relativity of acceleration in the Newtonian laws [Norton].
The fact that Newtonian laws are recovered when the mass of the universe tends to zero ($M_{u}\rightarrow 0\,,a_{0}\rightarrow 0$) fixing the gravitational constant, and that the MOND potential diverges at infinity, point towards the long-discussed problem of boundary conditions of the universe at large in GR, which has been linked to Mach’s principle in the past (and Einstein’s attempt to specify a divergence of the metric at infinity to satisfy the relativity of inertia).
But the mass of the universe can never be zero in the system under study, it can only be as small as the very same mass of the system itself. Decreasing the mass of the universe until the absence of any background ($M_{u}\rightarrow M,R_{u}\rightarrow r$), the function (4) reduces to some simple number $n$ which can be set to unity with the appropriate choice of integers in the formulation of the Machian interpolating function and in Sciama’s relationship, and velocities approach $v\rightarrow nc$. A simple way to obtain $n=1$ is considering Schrödinger’s or Treder’s definition of $G=c^{2}/(2M_{u}/R_{u})$ and $a_{0}=3GM_{u}/{R_{u}}^{2}$ from the scalar sum of $1/r^{2}$ masses for a uniform matter density sphere of total mass $M_{u}$. This is already a Machian feature of Sciama’s relationship, by which in absence of a background universe, rotation is undefined and tangential speed can take any value up to the limit of the speed of light. The proposed transformation respects this result as well as the invariance under global rescalings of mass, length, and time that Sciama’s relationship introduces in the Newtonian regime. Machian MOND extends such invariance to the deep MOND regime.
@@
Machian MOND satisfies several definitions of Mach’s principle: Newton’s gravitational constant $G$ is a dynamical field, and consequently, the inertial mass of a body increases with the agglomeration of masses in its neighborhood. Moreover, there’s an extra decrease of inertia a body is isolated from other masses (at low local gravitational field intensities) due to the interpolating function. The mass and size of the observable universe are always accounted for when considering non-inertial reference frames and local inertial frames are affected by the cosmic distribution of matter. Rotation is undefined up to the speed of light in the absence of the rest of the universe (or equivalently, a body in an empty universe has no inertia).
Milgrom himself favors modifying inertia (which is the true aim of Mach’s principle), and not just gravity. As modified inertia, the inertial law in Machian MOND depends on the inertial mass of the body, the gravitationally active mass $M$ and the mass of the universe $M_{u}$, i.e., inertia depends on all mass distributions at play, in a Machian spirit. Once inertia is understood as having an origin in gravitation under the Machian reasoning, the issues of MOND as modified inertia by a kinetic term in function of the trajectory (such as non-localities due to the function requiring knowledge of the entire trajectory [Milgrom1993]) no longer occur. Scalar matter fields, either inverse or inverse-squared distance ones, can be introduced in Newton’s second law, without modifying the Poisson equation (and leaving gravity intact).
Milgrom’s reluctance to explore a Machian version of MOND could be attributed to the widespread false claim that Mach’s principle necessarily implies some sort of anisotropy of inertia. The anisotropy of inertial mass is strongly constrained by the Hughes-Drever experiments, and nearby nonuniform mass distributions, such as the Milky Way in experiments on Earth, do not show a preferred direction for inertia. Not only was it originally pointed out as a possible but not necessary consequence of Mach’s principle [Cocconi1958], but the debate extended with Robert Dicke’s claims about the unobservability of anisotropic inertial mass effects mass in relativistic theories and Treder’s defense of scalar inertial mass in classical models, which prevents it from having a preferred direction.
@@
MOND’s interpolating function resembles the structure of a Lorentz-like transformation $\gamma=1/\sqrt{1-\beta}$. If a function depending on velocities, such as the Lorentz factor, arises from Lorentz invariance in SR, where accelerations are absolute, one would expect a function depending on accelerations or field intensities, such as Machian MOND, to arise when relativizing accelerations. The introduction of the constant speed of light from Maxwell’s electromagnetism into relativity lead to Einstein’s SR, and perhaps Milgrom’s acceleration scale constant inspires a new relativity of accelerations. Lorentz invariance implies that the speed of light is constant in all inertial frames, and it seems that Machian MOND implies that the universal field scale at any point in the universe is always $M_{u}/R_{u}^{2}$. It may be that Machian MOND arises from a new boost symmetry generalizing Lorentz invariance or from a transformation between accelerating frames, for instance, through the relativity of acceleration or the relativity of inertia, which is broken by the presence of the universe at large. Without a background universe, this symmetry would reduce to the Galilean or Lorentzian one.
@@
Machian MOND, or a similar and equivalent approximate formulation, is considered to be a necessary non-linear feature that a phenomenological theory of modified inertia or modified gravity which incorporates Mach’s principle reduces to as an approximation, in order to agree with galaxy rotation curves. This might arise by imposing Mach’s principle (for instance, through the relativity of inertia of local objects, being determined by the sources of the gravitational field at the large scale) to a non-Machian non-linear theory such as GR (in which solutions to local dynamics do not depend on the mass or size of the observable universe). In fact, analogous interpolating functions in SR and GR have, as independent variables, dimensionless ratios of the form $v/c$ and $GM/Rc^{2}$, similar to Machian MOND. The form of the function (4) is not unique and depends on the chosen Milgrom’s MOND interpolating function for which the Machian substitutions are made. Although (4) might be in conflict with solar system constraints, as it approaches $\mu(x)\rightarrow 1$ too slowly at the high acceleration regime, this is a problem of the slope of the function, which can be chosen to be different by another form for it. The function should arise naturally as an approximation, without the need of introducing it by hand (as introduced unnaturally in AQUAL or QUMOND), but the global $M_{u}$ and $R_{u}$ parameters governing local dynamics should appear in the complete FundaMOND theory, which is not the case in GR. So far, only linear and non-Lorentz invariant Machian theories of modified inertia with Sciama’s relationship have been constructed, such as those of Treder and Schrödinger.
@@
It is well known that, according to Bekenstein, a higher value for $a_{0}$ could resolve the tension between Milgrom’s MOND and galaxy cluster dynamics (and that $a_{0}$ seems to grow with scale), where Newtonian accelerations take around that same value (provided that $a_{0}$ stays the same in galaxies for agreement with rotation curves). Machian MOND is a varying $a_{0}$ model with external mass distributions and could in principle help resolving this issue of Milgrom’s MOND. This possibility will be discussed in a subsequent paper in preparation [Santander]. It could also solve the cosmological issues of Milgrom’s MOND [Felten], such as those arising from a fixed scale-dependence on $a_{0}$. As modified gravity, the equivalent variation of the Newtonian $G$ is both temporal due to the expanding universe and dependence on the mass and radius of the observable universe, and spatial since local potentials and local field intensities should also be included in Sciama’s relationship and the Machian MOND function, respectively.
@@
It would be interesting to compare Machian MOND with observational constraints of a varying effective $G$ and $a_{0}$ at the early universe, considering a cosmological model without most or all physical dark matter (assuming that MOND solves at least partly the need for dark matter), due to their dependence on the mass and radius of the observable universe, which have both varied over time. A similar study has been done for quantized inertia by showing the prediction of a specific increase in the galaxy rotation anomaly at higher redshifts [McCulloch2017], but quantized inertia only makes the $a_{0}$ of Milgrom’s MOND dependent on $R_{u}$ through $a_{o}\sim c^{2}/R_{u}$. The formulation of quantized inertia is similar to the one proposed (without Sciama’s relationship), but its interpretation is not based on Mach’s principle. In fact, no Planck constant appears in the formulation, which suggests that its origin is not related to quantum phenomena, while the need to account for the universe at large directly points towards Mach’s principle. More recently, it has been observed that there is no significant evolution in the TF law for redshifts up to $z\sim 2$ [Stacy], implying an almost constant $a_{o}$ in MOND. In contrast, Machian MOND shows a different dependence compared to quantized inertia: the Machian MOND function depends not only on $R_{u}$, but also on $M_{u}$. If the term $M_{u}/R_{u}^{2}$ is more constant than the simple dependence on $R_{u}$, this observation would not rule out Machian MOND (the observable radius was smaller in the past, but so was the mass contained in the volume defined by the observable radius). But as explained before, Machian MOND depends on many uncertainties, such as the Hubble tension, whether some physical dark matter is present in $M_{u}$, missing integers and Pi in the formulation, and relativistic effects of the universe as a whole.
@@ Acknowledgments
The authors have no competing interests to declare. This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.